### v1

@@ Results
To assess the diversity of the user bases across platforms, we evaluate the distribution of users’ political leaning. We infer the political leaning of active users (i.e., users who shared 10 or more URLs toward domains with an associated political bias) based on their posting activity. We assign a numerical score to news media, ranging from $-1$ (extreme-left) to $+1$ (extreme-right), excluding unreported sources. The political leaning of each user is then defined as the average of the leanings of the news domains they post (see section Materials and Methods for further details).
Fig. 4 illustrates the political leaning distribution of users in the nine platforms. We note that the user bases of Facebook and Twitter are composed of two distinct groups spanning the whole ideological spectrum, resulting in a bimodal distribution of political leanings. This result suggests the presence of segregated communities endorsing opposite narratives. The same pattern is not observed on some alt-tech platforms (Gab, Parler, and Scored), where users are concentrated within a narrower leaning interval, forming a homogeneous, right-leaning user base. Reddit exhibits similar behavior to alt-tech platforms but with a homogeneous user base skewed significantly leftward, in line with previous findings [46, 47]. Voat and BitChute stand out among the alt-tech platforms, displaying a relatively heterogeneous user base compared to their counterparts. For Voat, this heterogeneity likely stems from the significant portion of traffic redirected toward left-biased sources by alt-right groups, as previously observed. For BitChute, the high proportion of unreported sources on the platform likely contributes to its heterogeneity. Specifically, since we only include users who shared at least 10 URLs with an associated political bias in our computation, this criterion captures only a small, seemingly diverse subset of the user base.
@@
Figure 4: Distributions of users’ political leaning for each platform. Each unique user gets assigned a leaning score between $-1$ and $+1$, according to their posting activity. The bars are colored according to the number of “questionable” sources shared by users of a specific leaning. We notice how Facebook and Twitter have a polarized user base with two distinct groups, one sharing mostly reliable content and the other sharing mostly content with low factual reporting. The situation is more homogeneous regarding all of the fringe platforms, and the remaining two mainstream platforms.
@@
Additionally, on BitChute, a substantial share of content from hyper-extremist sites is driven by a very small group of highly active users. In some cases, these users are only a few dozen or hundreds, yet they are responsible for thousands of links, creating a disproportionate volume of extremist content on the platform. This finding aligns with the concept of “vocal minorities” [48], where a small but highly active subset of users exerts an outsized influence on the platform by driving a significant share of its activity. Interestingly, users with a political leaning below 0 (left-leaning) make up approximately 30% of the total user base but share, on average, 130 fewer links than right-leaning users (leaning above 0) and 217 fewer links than those with a leaning equal or above $0.5$. This suggests that left-leaning users are generally less active in sharing content, resulting in a smaller contribution to the circulation of news on the platform. These dynamics underscore an asymmetry in content production and dissemination across ideological groups, further highlighting the influence of vocal minorities in shaping platform-wide narratives.
@@
The observations regarding the diversity of the user bases can be quantified by means of the variance of the political leaning distribution, $\sigma^{2}$, reported in Table 1. Platforms such as Twitter, Facebook, YouTube, and BitChute exhibit high variance, indicating a wide spread of user leanings and reflecting a diverse user base, while the remaining platforms demonstrate more homogeneous user bases.
When considering the reliability of the sources shared, we observe that users with right-leaning and extreme-right political orientations share a higher proportion of questionable content. However, this pattern aligns with established correlations reported by MBFC and is not specific to our study. Therefore, we focus on broader patterns of user base composition and political leanings rather than the relationship between political orientation and source reliability.
@@ Conclusions
This study introduced the concept of “echo platforms”, i.e., entire social media platforms that operate as self-contained echo chambers, reinforcing homogeneous beliefs and isolating users from opposing viewpoints. Echo platforms represent an evolution of traditional echo chambers, where the platform structure fosters ideological uniformity. Our framework categorizes platforms across three dimensions—platform centrality, content reliability, and user base homogeneity—providing a systematic approach to understanding these dynamics within the broader context of internet fragmentation. Analyzing a multi-platform dataset spanning mainstream social media (e.g., Facebook, Twitter) as well as alt-tech platforms (e.g., Gab, Parler), we revealed patterns of content redirection, community insulation, and polarized news consumption. This systemic shift from segregated communities to platform-wide segregation highlights the role of moderation policies, user self-selection, and engagement-driven business models in shaping the rise of echo platforms.
@@
Our findings contribute to the discourse on digital polarization, showing how alt-tech platforms attract marginalized or de-platformed communities from mainstream spaces, creating distinct niches that intensify ideological separation. This underscores the need to reevaluate moderation strategies, as current policies may unintentionally exacerbate the fragmentation of the information ecosystem by fostering the growth of insulated echo platforms. Such dynamics challenge the digital public sphere by limiting constructive dialogue and deepening ideological divides.
@@
While our operational framework provides robust tools for identifying and analyzing echo platforms, the study has limitations. The dataset, though extensive, reflects a specific period marked by significant political events, such as elections. This context offers valuable insights into platform behavior during critical societal moments but may only partially capture ongoing changes in the social media landscape. Future research could extend this work by integrating more recent data and exploring cross-cultural or longitudinal variations in platform dynamics.
@@
Our results offer a foundation for examining the socio-political impacts of echo platforms, particularly in contexts where public discourse and social cohesion are at risk. Understanding and mitigating the societal effects of echo platforms is essential for developing policies that balance open expression with fostering a diverse and cohesive public sphere. This study underscores the critical role of echo platforms in shaping the modern information ecosystem, emphasizing their contribution to internet fragmentation and societal polarization. By offering a detailed quantitative framework, we aim to advance understanding of these platforms and inform strategies to mitigate the adverse impacts of digital fragmentation on society.
Matherial and Methods
@@
In this section, we describe the procedures for data collection and preprocessing, along with details about the datasets used in our study. During data preprocessing, we manually excluded URLs that were not relevant to our analysis. This included self-links (i.e., a platform linking to itself), URLs directing to social media platforms not included in the datasets (e.g., Snapchat, TikTok), video streaming services (e.g., Vimeo, Dailymotion), financial services (e.g., PayPal, Venmo), music streaming services (e.g., Spotify, SoundCloud), and tech platforms or services (e.g., Google Suite, Streamlabs). Additionally, we removed a range of URLs that did not fit these broad categories but were similarly uninformative for our purposes, such as links to Amazon, Steam, Coinbase, and NASA.
@@
Twitter: Our dataset is based on information retrieved from Flamino et al. [43] and includes 174 million tweets posted between June 1 and November 3, 2020.
@@
YouTube: We use URLs extracted from the descriptions of 270,000 YouTube videos, collected using the YouTube Data API between June and December 2020. The videos were identified through searches based on predefined keyword lists: L1: Trump, trump, #donaldtrump, #trump and L2: Biden, biden, #joebiden, #biden. For each list, an additional search was conducted by crawling the network of videos using YouTube’s algorithm to identify related videos*. From the gathered dataset, we retained only videos containing $Trump$$trump$ (or $Biden$$biden$, respectively) in the title or description.
@@
BitChute: BitChute is a British alt-tech video hosting platform known for its lower moderation efforts compared to its more popular counterpart, YouTube. The platform focuses heavily on news and politics and is associated with a significant amount of hate speech in both videos and comment sections [50]. For our analysis, we use the MeLa BitChute Dataset [51], which provides a near-complete scrape of the platform from 2019 to 2021. From this dataset, we filter the data to retain only the 1.1 million videos uploaded in 2020, out of the original 3.3 million videos.
@@
Gab: Gab is an alt-tech microblogging platform structured similarly to Twitter but with minimal content moderation, promoting free speech and Christian values [52, 53]. With its predominantly far-right user base, Gab has been described as a safe haven for neo-Nazis, members of the American alt-right, Trump supporters, and conspiracy theorists [54, 55]. The platform has also been repeatedly linked to online radicalization and real-world violent events [56].
@@
The data for this study was collected in two phases. From June 1, 2020, to October 23, 2020, posts were downloaded using Gab’s general stream, capturing all posts generated during this period. After October 23, due to the deprecation of the API endpoint, data was collected using the timelines of 930,000 users identified in the first phase. The dataset was then filtered using the keywords trump” or biden,” resulting in 467,000 posts made between June 1 and December 1, 2020.
@@
Parler: Launched in 2018, Parler is a platform with functionalities similar to Twitter, marketed as a free-speech-focused alternative. It gained mainstream attention following the storming of Capitol Hill in 2021, as it was one of the platforms allegedly used to incite and plan the attack [57, 58]. Shortly after, Parler was removed from the Google Play Store and Apple App Store and subsequently suspended by its hosting provider, Amazon AWS. At the time of writing, Parler’s social media outlet has announced plans to relaunch its services, though the exact timeline remains unclear.
@@
The data used in this study comes from a large dataset comprising 183 million posts made between 2018 and 2021, collected by Aliapoulios et al. [59]. For our analysis, we focused exclusively on the 13 million posts made during 2020.
@@
Scored: Scored (accessible via both https://scored.co and https://communities.win) emerged as an alternative to Reddit, hosting numerous communities that were banned from more prominent social networks, including $c/TheDonald$, $c/GreatAwakening$, and $c/FatPeopleHate$. For this study, we use the iDRAMA-Scored-2024 dataset [60], a comprehensive scrape of the platform since its inception in 2020. This dataset includes a total of 6.2 million posts spanning nearly four years, of which 1.6 million posts from 2020 are analyzed in our work.
@@
Voat: The now-defunct Voat, which was shut down in December 2020, served as an alternative to Reddit, similar to Scored. Several communities banned from Reddit migrated to Voat, forming new subverses—the platform’s equivalent of subreddits. Notable examples include v/fatpeoplehate, v/TheRedPill, and v/GreatAwakening. For this study, we use data from a collection of 2.3 million posts made on the platform between November 2013 and December 2020, compiled by Mekacher and Papasavva [61]. From this dataset, we retain only submissions made in 2020, excluding those posted in the subverse v/QRV due to its policy of anonymizing user IDs, resulting in a total of 355,000 submissions.
@@
The set of connections (URLs) between platforms can be modeled as a weighted directed graph $G=(V,E,w)$, where $V$ represents the set of nine social media platforms, $E\subseteq V\times V$ the set of directed edges between platforms, and $w:E\to\mathbb{R}^{+}$ the weight of edges corresponding to the number of times one platform links to another. Given the size disparity between the considered data sets, quantifying the relevance of links among nodes using only their weights would introduce a bias. Thus, we compute how such weights deviate from their expected values. These expected values can be computed using the weighted configuration model [62], a null model in which the in- and out-strength distributions of the nodes are kept. According to the weighted configuration model, the average weight of a link connecting two uncorrelated vertices with out-strength $s^{out}_{i}$ and in-strength $s^{in}_{j}$ can be written as $\mathbb{E}[w_{ij}]=\frac{{s^{out}_{i}\cdot s^{in}_{j}}}{S}$, where $S$ refers to the total weight of the network. Hence, we can compute the relationship between the observed weights and their expectation as follows:
@@
As mentioned, we utilize Media Bias/Fact Check (MBFC, https://mediabiasfactcheck.com) to label news outlets based on their political bias and level of reliability regarding factuality reporting. MBFC is an independent fact-checking organization that rates various news sources. The labeling utilized in this study, collected in October 2024, contains political bias categories ranging from extreme-left to extreme-right, while certain news outlets are classified as “questionable”, indicating that they exhibit one or more of the following, per MBFC: “extreme bias, consistent promotion of propaganda/conspiracies, poor or no sourcing to credible information, a complete lack of transparency and/or is fake news”. While not explicitly falling under the “questionable” definition of MBFC, we also consider as such domains classified as “conspiracy/pseudoscience”, given their inherent low credibility and absence of proper fact-checking. Furthermore, we manually labeled a small number of right-biased and/or extremist websites which accounted for a significant portion of some alt-tech platforms’ traffic, but that were not present on MBFC.
@@
Table 3: Timeframe and number of URLs collected (before and after processing) for each platform’s data set.
@@
Parler 01/01/2020 - 31/12/2020 14M 8.2M
Scored 01/01/2020 - 31/12/2020 984k 883k
Voat 01/01/2020 - 25/12/2020 269k 215k
Table 4: Table showing the name, number of URLs shared, number of URLs shared toward left biased sources, fraction of left biased sources shared, cumulative count and cumulative proportion of left biased sources shared for the eleven Voat’s subverses responsible for sharing approximately 80% of the platform’s left biased content. Note how in this table we consider as “left biased” every source classified as either extreme-left biased, left biased, or left-center biased.
@@
1. AnonAll 15822 9643 $0.609$ 9643 $0.298$
2. news 31884 5931 $0.186$ 15574 $0.482$
3. politics 20898 2627 $0.126$ 18201 $0.563$
4. whatever 25841 2327 $0.090$ 20528 $0.635$
5. GreatAwakening 35987 1535 $0.043$ 22063 $0.682$
6. OccidentalEnclave 4865 1345 $0.276$ 23408 $0.724$
7. Niggers 5621 553 $0.098$ 23961 $0.741$
8. Worldnews 2289 550 $0.240$ 24511 $0.758$
9. theawakening 12765 446 $0.035$ 24957 $0.772$
10. technology 2780 442 $0.159$ 25399 $0.785$
11. Australia 919 388 $0.422$ 25787 $0.797$
Figure 5: Adjacency matrix describing the number of links from and to each of the nine different platforms.
Figure 6: Matrix showing the weighted cosine similarity between the top-20 linked domains for each platform.

### v2

@@ Results
To assess the diversity of the user bases across platforms, we evaluate the distribution of users’ political leaning. We infer the political leaning of active users (i.e., users who shared 10 or more URLs toward domains with an associated political bias) based on their posting activity. We assign a numerical score to news media, ranging from $-1$ (extreme-left) to $+1$ (extreme-right), excluding unreported sources. The political leaning of each user is then defined as the average of the leanings of the news domains they post (see section Materials and Methods for further details). We note that YouTube and BitChute operate on different principles compared to the other platforms considered. As video-sharing platforms, they lack traditional feeds or direct user interactions. Consequently, our estimation of political leanings on these platforms reflects the tendencies of active content producers rather than user consumption patterns. Additionally, we acknowledge that using the average as a summary statistic may fail to capture behaviors more evenly distributed across a bias spectrum: as such, a robustness check showing the distribution of the variance of the political leaning scores assigned to users is available in SI (Fig. S4). The analysis of the variance at the user level shows that most of them tend to engage with content from a narrow ideological range, making the average a quite representative summary statistic of the general trend.
@@
Figure 4: User base composition. Distributions of users’ political leaning for each platform. Each unique user gets assigned a leaning score between $-1$ and $+1$, according to their posting activity. The bars are colored according to the number of “questionable” sources shared by users of a specific leaning. We notice how Facebook and Twitter have a polarized user base with two distinct groups, one sharing mostly reliable content and the other sharing mostly content with low factual reporting. The situation is more homogeneous regarding all of the fringe platforms, and the remaining two mainstream platforms.
Fig. 4 illustrates the political leaning distribution of users in the nine platforms. We note that the user bases of Facebook and Twitter are made up of two distinct groups that span the entire ideological spectrum, resulting in a bimodal distribution of political leaning. The same pattern is not observed on some alt-tech platforms (Gab, Parler, and Scored), where users are concentrated within a narrower leaning interval, forming a homogeneous, right-leaning user base. Interestingly, Reddit exhibits a skewed leaning distribution, similar to alt-tech platforms, but concentrated on left leaning content, as noted in previous work [51, 52]. Voat and BitChute display a slightly more heterogeneous user base compared to other alt-tech platforms. For Voat, this heterogeneity likely stems from the significant portion of traffic redirected toward left-biased sources by alt-right groups, as previously observed. For BitChute, the high proportion of unreported sources on the platform likely contributes to its heterogeneity. Specifically, since we only include users who shared at least 10 URLs with an associated political bias in our computation, this criterion captures only a small subset of the user base.
@@
Expanding on BitChute, users with a political leaning below 0 (left-leaning) make up approximately 12% of the total user base but share, on average, 15 fewer links than right-leaning users (leaning above 0) and 30 fewer links than those with a leaning equal or above $0.5$. This suggests that left-leaning users are generally less active in sharing content, resulting in a smaller contribution to the circulation of news on the platform. These dynamics underscore an asymmetry in content production and dissemination across ideological groups, highlighting the influence of vocal communities in shaping platform-wide narratives.
@@
The observations regarding the diversity of the user bases can be quantified by means of the variance of the political leaning distribution, $\sigma^{2}$, reported in Table 1. Platforms such as Twitter, Facebook, YouTube, BitChute, and Voat to some extent, exhibit higher variance, indicating wide spread user leanings and reflecting a diverse user base, while the remaining platforms demonstrate more homogeneous user bases. When considering the reliability of the sources shared, we observe that users with right-leaning and extreme-right political orientations share a higher proportion of questionable content. However, this pattern aligns with established correlations reported by MBFC and is not specific to our study.
@@ Conclusions
This study characterized fragmentation in the social media ecosystem related to the 2020 US presidential election, and introduced the concept of “echo platforms”, i.e., entire social media platforms that operate as self-contained echo chambers, reinforcing homogeneous beliefs and isolating users from opposing viewpoints. Echo platforms represent an evolution of traditional echo chambers, where the platform structure fosters ideological uniformity. Our framework categorizes platforms across three dimensions—platform centrality, content reliability, and user base homogeneity—providing a systematic approach to understanding these dynamics within the broader context of social media fragmentation.
Such analysis provides a better characterization of the differences between mainstream platforms (e.g., Facebook, Twitter) and alt-tech ones (e.g., Gab, Parler). In particular, we uncover that Reddit shares some features with mainstream platforms, like the high reliability of posted news media, but others with fringe platforms, such as a uniform user composition and a relatively peripheral role in cross-platform linking. Our findings suggest that, despite being a mainstream social media platform, Reddit functions more like an echo-platform than others.
@@
This systemic shift from segregated communities to platform-wide segregation highlights the role of moderation policies, user self-selection, and engagement-driven business models in shaping the rise of echo platforms. Our findings contribute to shedding light on the factors influencing digital polarization, showing how alt-tech platforms attract marginalized or de-platformed communities from mainstream spaces, creating distinct niches that intensify ideological separation. This underscores the need to reevaluate moderation strategies, as current policies may unintentionally exacerbate the fragmentation of the social media ecosystem by fostering the growth of insulated echo platforms. Such dynamics challenge the digital public sphere by limiting constructive dialogue and deepening ideological divides.
@@
While our framework provides robust tools for identifying and analyzing echo platforms, the study has limitations. The dataset, though extensive, reflects a specific period marked by significant political events, such as elections. This context offers valuable insights into platform behavior during critical societal moments but may only partially capture ongoing changes in the social media landscape. Future research could extend this work by integrating more recent data and exploring cross-cultural or longitudinal variations in platform dynamics, as well as investigating the impact of coordinated actions and in the fragmentation of online spaces, and, where feasible, incorporating cross-platform user identification to enhance the depth of analysis.
@@
Our results offer a foundation for examining the socio-political impacts of echo platforms, particularly in contexts where public discourse and social cohesion are at risk. Understanding and mitigating the societal effects of echo platforms is essential for developing policies that balance open expression with fostering a diverse and cohesive public sphere.
This study highlights the influence of echo platforms in shaping today’s social media landscape, particularly their role in internet fragmentation and societal polarization. By offering a detailed quantitative approach, we aim to advance understanding of these platforms and inform strategies to mitigate the adverse impacts of digital fragmentation on society.
Methods
@@
In this section, we describe the procedures for data collection and preprocessing, along with details about the datasets used in our study. All of the datasets were filtered to retrieve politically relevant content pertaining to the 2020 U.S. election, either by performing a keyword search, or by selecting relevant political communities where a keyword search was not feasible. The set of keywords is consistent among the datasets in which they were utilized (with minor variations), and based on the presidential candidates’ names. During data preprocessing, we manually excluded URLs that were not relevant to our analysis. This included self-links (i.e., a platform linking to itself), URLs directing to social media platforms not included in the datasets (e.g., Snapchat, TikTok), video streaming services (e.g., Vimeo, Dailymotion), financial services (e.g., PayPal, Venmo), music streaming services (e.g., Spotify, SoundCloud), and tech platforms or services (e.g., Google Suite, Streamlabs). Additionally, we removed a range of URLs that did not fit these broad categories but were similarly uninformative for our purposes, such as links to Amazon, Steam, Coinbase, and NASA.
@@
Twitter: The dataset [54], based on information retrieved from Flamino et al. [46], was obtained using the Twitter Search API, with the names of the two U.S. presidential candidates from the 2020 election as keywords (consistently with Facebook). It includes 174 million tweets posted between June and November 2020.
@@
YouTube: We use URLs extracted from the descriptions of 270,000 YouTube videos, collected using the YouTube Data API between June and December 2020. The videos were identified through searches based on the same keywords utilized for Facebook. For each video, an additional search was conducted by crawling the network of related videos, as suggested by YouTube’s algorithm. From the gathered dataset, we retained only videos containing $Trump$$trump$ (or $Biden$$biden$, respectively) in the title or description.
@@
BitChute: BitChute is a British alt-tech video hosting platform known for its lower moderation efforts compared to its more popular counterpart, YouTube. The platform focuses heavily on news and politics and is associated with a significant amount of hate speech in both videos and comment sections [55]. For our analysis, we use the MeLa BitChute Dataset [56], which provides a near-complete scrape of the platform from 2019 to 2021, from which we filtered the data to retain only the 1.1 million videos uploaded in 2020. We further refined the dataset by performing a keyword search on the video titles. The keywords used for filtering were {trump, biden, joebiden, donaldtrump, donald, trump2020, biden2020}, resulting in the retention of 40 thousand videos.
@@
Gab: Gab is an alt-tech microblogging platform structured similarly to Twitter but with minimal content moderation, promoting free speech and Christian values [57, 58]. With its predominantly far-right user base, Gab has been described as a safe haven for neo-Nazis, members of the American alt-right, Trump supporters, and conspiracy theorists [59, 60]. The platform has also been repeatedly linked to online radicalization and real-world violent events [61, 37].
@@
The data for this study was collected in two phases. From June 1, 2020, to October 23, 2020, posts were downloaded using Gab’s general stream, capturing all posts generated during this period. After October 23, due to the deprecation of the API endpoint, data was collected using the timelines of 930,000 users identified in the first phase. The dataset was then filtered using the keywords {“trump”, “biden”}, resulting in 467,000 posts made between June 1 and December 1, 2020.
@@
Parler: Launched in 2018, Parler is a platform with functionalities similar to Twitter, marketed as a free-speech-focused alternative. It gained mainstream attention following the storming of Capitol Hill in 2021, as it was one of the platforms allegedly used to incite and plan the attack [35, 36]. Shortly after, Parler was removed from the Google Play Store and Apple App Store and subsequently suspended by its hosting provider, Amazon AWS. At the time of writing, plans to relaunch Parler’s services have been announced, though the exact timeline remains unclear.
@@
The data used in this study comes from a large dataset comprising 183 million posts made between 2018 and 2021, collected by Aliapoulios et al. [62]. For our analysis, we focused exclusively on the 1.7 million posts made during 2020 containing at least one of the keywords {trump, biden, joebiden, donaldtrump, donald, trump2020, biden2020} in their main body.
@@
Scored: Scored (accessible via both https://scored.co and https://communities.win) emerged as an alternative to Reddit, hosting numerous communities that were banned from more prominent social networks, including $c/TheDonald$, $c/GreatAwakening$, and $c/FatPeopleHate$. For this study, we use the iDRAMA-Scored-2024 dataset [63], a comprehensive scrape of the platform since its inception in 2020, for a total of 6.2 million posts spanning nearly four years. We filter the dataset by retaining only the submissions posted during 2020 in six politically centered communities of the 975 present communities, which account for 92% of links shared during 2020 over 1.4M posts. Please refer to the SI for a detailed table of the selected communities retained in the analysis (Table S3).
@@
Voat: The now-defunct Voat, which was shut down in December 2020, served as an alternative to Reddit, similar to Scored. Several communities banned from Reddit migrated to Voat, forming new subverses—the platform’s equivalent of subreddits. Notable examples include v/fatpeoplehate, v/TheRedPill, and v/GreatAwakening. For this study, we use data from a collection of 2.3 million posts made on the platform between November 2013 and December 2020, compiled by Mekacher and Papasavva [64]. From this dataset, we retain only submissions made during 2020 across the 60 most prominent political subvoats (over the $7,604$ total subvoats), for a total of almost 200 thousands posts. Please refer to the SI for a table of the selected subverses retained in our analysis (Table S4).
@@
The set of connections (URLs) between platforms can be modeled as a weighted directed graph $G=(V,E,w)$, where $V$ represents the set of nine social media platforms, $E\subseteq V\times V$ the set of directed edges between platforms, and $w:E\to\mathbb{R}^{+}$ the weight of edges corresponding to the number of times one platform links to another. Given the size disparity between the considered data sets, quantifying the relevance of links among nodes using only their weights would introduce a bias. Thus, we compute how such weights deviate from their expected values. These expected values can be computed using the weighted configuration model [65], a null model in which the in- and out-strength distributions of the nodes are preserved. According to the weighted configuration model, the average weight of a link connecting two uncorrelated vertices with out-strength $s^{out}_{i}$ and in-strength $s^{in}_{j}$ can be written as $\mathbb{E}[w_{ij}]=\frac{{s^{out}_{i}\cdot s^{in}_{j}}}{S}$, where $S$ refers to the total weight of the network. Hence, we can compute the relationship between the observed weights and their expectation as follows:
@@
As mentioned, we utilize Media Bias/Fact Check (MBFC, https://mediabiasfactcheck.com/) to label news outlets based on their political bias and level of reliability regarding factuality reporting. MBFC is an independent fact-checking organization that rates various news sources. The labeling utilized in this study, collected in October 2024, contains political bias categories ranging from extreme-left to extreme-right, while certain news outlets are classified as “questionable”, indicating that they exhibit one or more of the following, per MBFC: “extreme bias, consistent promotion of propaganda/conspiracies, poor or no sourcing to credible information, a complete lack of transparency and/or is fake news”. While not explicitly falling under the “questionable” definition of MBFC, we also consider as such domains classified as “conspiracy/pseudoscience”, given their inherent low credibility and absence of proper fact-checking. Furthermore, we manually labeled a small number of right-biased and/or extremist websites which accounted for a significant portion of some alt-tech platforms’ traffic, but that were not present on MBFC.
@@
Table S1: Timeframe and number of URLs collected (before and after processing) for each platform’s data set.
@@
Parler 01/01/2020 - 31/12/2020 14M 1.1M
Scored 01/01/2020 - 31/12/2020 984k 868k
Voat 01/01/2020 - 25/12/2020 269k 135k
Table S2: Table showing the subverse name, number of URLs shared, number of URLs shared toward left biased sources, fraction of left biased sources shared, cumulative count and cumulative proportion of left biased sources shared for the six Voat’s subverses responsible for sharing 89.5% of the platform’s left biased content. Note how in this table we consider as “left biased” every source classified as either extreme-left biased, left biased, or left-center biased.
@@
1. AnonAll 15822 9643 $0.609$ 9643 $0.399$
2. news 31884 5931 $0.186$ 15574 $0.644$
3. politics 20898 2627 $0.125$ 18201 $0.753$
4. GreatAwakening 35987 1535 $0.042$ 19736 $0.816$
5. OccidentalEnclave 4865 1345 $0.276$ 21081 $0.872$
6. Niggers 5621 553 $0.098$ 21634 $0.895$
Table S3: Selected Scored Communities retained in our analysis, and number of URLs contained in them
# Community Nr. of URLs
1 TheDonald 907439
2 OmegaCanada 7034
3 GreatAwakening 2804
4 GavinMcInnes 1248
5 Conspiracies 1180
6 Conservative 224
Table S4: Selected Subverses retained in our analysis, and number of URLs contained in them.
# Subverse name Nr. of URLs
1 GreatAwakening 35987
2 news 31884
3 politics 20898
4 AnonAll 15822
5 theawakening 12765
6 Niggers 5621
7 OccidentalEnclave 4865
8 Conspiracy 3766
9 WorldToday 2753
10 Worldnews 2289
11 TheDonald 1620
12 TheGreatAwakening 920
13 USNews 776
14 politicalnews 754
15 2020ElectionNews 749
16 Worldpolitics 693
17 uspolitics 451
18 CRIMENEWS 431
19 Jews 359
20 christiannews 284
21 ChristianEnclave 281
22 CultureWars 278
23 economics 250
24 SJWHate 221
25 PivottoAsia 219
26 AnonNews 215
27 Libertarian 161
28 pizzagate 160
29 CanadaFirst 150
30 Military 145
31 pedogate 140
32 PoliticalDiscussion 123
33 BadCopNoDonut 112
34 pizzagateuncensored 112
35 education 111
36 HillaryforJail 76
37 Lawenforcement 73
38 PoliticallyIncorrect 69
39 BlackLivesMatter 68
40 conspiracyfact 66
41 newsandpolitics 61
42 PedogateFullExposure 59
43 WhiteRights 58
44 ANTIFAWATCH 55
45 gunpolitics 54
46 Immigration 54
47 Identitarians 47
48 ClimateSkeptics 44
49 ObamaForPrison 44
50 QAnon 44
51 Censorship 43
52 CrimesByRace 40
53 Jewspiracy 40
54 russia 40
55 War 39
56 Kikes 38
57 Abortion 37
58 MassDeport 37
59 Israel 36
60 presstitutes 36
Table S5: Percentage of domains matched with MBFC ratings for each platform.
Platform % of rated domains
Facebook 5.85
Reddit 26.62
Twitter 3.94
YouTube 11.11
Bitchute 18.84
Gab 27.77
Parler 18.98
Scored 22.06
Voat 34.12
Figure S1: Adjacency matrix describing the number of links from and to each of the nine different platforms.
Figure S2: Cumulative fraction of the total amount of links towards different domains. As observed, the top-20 most shared domains encompass for all platforms a significant percentage of links shared.
Figure S3: Cosine similarity matrix for different thresholds (top 10, 20, 30, 50 domains). The results remain extremely similar between the different matrices.
Figure S4: Probability density function of the variance in the bias of the sources shared by users on different platforms. We observe that users on different platforms generally exhibit low variance in the political bias of the sources they share, suggesting that most tend to engage with content from a narrow ideological range.
Figure S5: Common time window of the data sets.
Figure S6: comparison between the rescaled adjacency matrix computed with the “full” datasets, and those obtained after filtering the datasets on the common time window. The results remain largely consistent between the two, both highlighting a fringe ecosystem generally more connected than the mainstream one. This similarity between the results is confirmed by computing Kendall’s Tau between the two matrices, which has a very high score of $\tau=0.92$.
Figure S7: Comparison of the original platform diet to the one computed after filtering the data sets maintaining the uniform time window. The results are practically the same, as also shown by the extremely high Kendall’s Tau value between the matrices of $\tau=0.967$.
Figure S8: Cosine similarity matrix computed on both the original and filtered dataframes. The values between the two are extremely similar, as also shown by the high value of Kendall’s Tau between the matrices ($\tau=0.966$). The only notable difference is an increase in the similarity between YouTube and Parler.
Figure S9: Overlap between original user leaning, and those obtained after filtering the data. We can observe that the histograms nearly completely overlap. This similarity is also shown by the reported values of the Jensen-Shannon divergence between the histograms, which results very low in every case considered.